# Towards Truly Zero-shot Compositional Visual Reasoning with LLMs as Programmers

### v1

@@ 1 Introduction
Compositional visual question answering requires a model to answer questions about visual content in a compositional manner, involving multiple steps of reasoning or considering relationships between different objects or entities within an image. It is a complex task as it requires understanding both the visual information in an image and the structure of the question, and reasoning about how different visual elements relate to one another to generate the correct answer. Recently, large progress has been made on many such vision and language tasks by scaling end-to-end neural networks models in terms of size, training data, and compute (Alayrac et al., 2022; Chen et al., 2022b; Yu et al., 2022; Wang et al., 2022a; Gan et al., 2022; Lu et al., 2022; Li et al., 2023; Driess et al., 2023; Chen et al., 2023d; c). However, even the largest state-of-the-art (SotA) models struggle in tasks that require compositional reasoning, ability to generalize, fine-grained spatial reasoning capabilities, and counting (Bugliarello et al., 2023; Paiss et al., 2023; Hsieh et al., 2023; Yuksekgonul et al., 2022; Zhao et al., 2022; Hendricks & Nematzadeh, 2021). An example of such task is the following query: “Could the cookies on the table be equally distributed among children?” (Surís et al., 2023). To solve this, the model needs to detect the cookies in the image, filter out the ones that are not on the table, detect children, count cookies and children, and check if the cookies count is divisible by the children count. Questions like these are difficult for current end-to-end vision and language models (VLMs). Scaling VLMs further makes them even more data- and compute-hungry(Villalobos et al., 2022), so the scale alone seems unlikely to solve these tasks, especially due to the exponentially-large long tail of compositional tasks.
@@
Recently, an alternative approach based on “tool use” gained popularity (Cobbe et al., 2021; Komeili et al., 2021; Thoppilan et al., 2022; Parisi et al., 2022; Zeng et al., 2022; Gao et al., 2023; Qin et al., 2023; Zhuge et al., 2023). In “tool use”, an LLM solves a task by controlling (akin to System 2) a set of tools (such as an object detector, akin to System 1) (Zeng et al., 2022; Shen et al., 2023; Zhuge et al., 2023). In particular, VisProg (Gupta & Kembhavi, 2023), ViperGPT (Surís et al., 2023), and CodeVQA (Subramanian et al., 2023) show great promise in solving visual question answering by using an LLM to generate a program (in Python or a custom scripting language). During execution, the program calls individual vision models (such as object detector, depth estimator) through an API that is provided in the prompt. For example, to answer “Which color is the jacket of the second person from the left?” (Figure 0(a)), the program needs to detect people, sort them from left to right, select the second, detect their jacket, and query its color. These models achieved SotA on compositional visual question answering, visual grounding, and video temporal reasoning tasks. By their construction, they are interpretable, compositional, adaptable (tools can be upgraded on the fly), offer strong generalization, mathematical, and reasoning skills, and do not require gradient-based training. However, in their current form, they heavily rely on human engineering of in-context (program) examples (ICEs) in the prompt. Moreover, ICEs are dataset- and task-specific. To generate them, significant labor by highly skilled programmers is required. For this reason, we argue that these methods should not be called zero-shot in their current form.
@@ 2.2 Abstract API through visual routines
When programming, we continuously build new layers of abstraction. We start from a set of primitive functions and then abstract them away by adding new functions. These are then grouped into libraries, and additional layers of abstraction are built on top. By abstracting away the implementation details, we are able to build systems with ever-increasing capabilities and complexity. Motivated by this, we introduce a set of spatially and temporally abstract functions (routines  [^1] Note that our “routines” do not correspond to the visual routines of Ullman (1987) such as tracing or scanning.) that abstract away a number of lines of code for the same functionality and together make the Abstract API. From a practical perspective, we are motivated by a number of failure cases observed in the experiments (see Section 3). As presented in Section 2.1, ViperGPT’s API is fairly simple (contains almost exclusively functions to call pretrained models). Although simplicity is good and often desirable, in the case of visual reasoning with LLMs as programmers, the lack of expressive visual routines requires the code-generating LLM to have strong spatial and temporal reasoning capabilities. Qualitative analysis showed that this is often not the case and that the current LLMs are not yet perfect in these terms (e.g. they confuse left and right, top and bottom). For example, for the query “return the second person from the right”, the program generated by the LLM correctly sorts the persons along the horizontal axis but then wrongly takes the second index in the array (instead of the second last). Similarly, they sometimes “confuse” temporal order, e.g., if a “before” event means a smaller or a larger time index.
@@ 2.3 Automatic generation of in-context examples
In-context examples (ICEs) greatly influence the performance of LLMs (Brown et al., 2020; Chen et al., 2023b). For example, ViperGPT (Surís et al., 2023) uses between 10 and 16 hand-engineered dataset-specific query-code ICEs per dataset. Similarly, VisProg (Gupta & Kembhavi, 2023) uses between 16 and 31 ICEs and CodeVQA (Subramanian et al., 2023) about 50 ICEs. However, constructing these ICEs requires heavy human engineering, as they might need to be rewritten in a way that the LLM can use them to “generalize” to other examples across the dataset. Furthermore, the constructed examples are specific not only to the dataset but also to the LLM and the API. If any of those changes, they need to be written from scratch. Finally, to write good query-code ICEs, highly skilled labor is required, ideally someone familiar with the workings of LLMs and a good grasp of Python programming.
@@ 2.4 Self-correction
In some cases, we know that code execution failed due to some components of a particular module. For example, the open vocabulary detector fails due to a too high threshold hyperparameter. When the threshold is high, we have a higher number of false negatives. For such cases, we propose to automatically change the hyperparameter of the module (e.g. reduce the threshold) and execute code again.
@@ 3 Experiments
For code generation, we use a code instruction-tuned version of PaLM 2 (Anil et al., 2023) code-bison accessible via the Google Cloud API (Google, 2023). We use the same model to select an answer for the multichoice questions in the NExT-QA dataset. Vision models we use are OWLv2 (Minderer et al., 2023) for object detection, SigLiT (Zhai et al., 2023) for text-image comparison, MiDaS (Ranftl et al., 2020) for depth estimation, and PaLI-3 (Chen et al., 2023d) for image captioning and answering visual queries. Note that all models are different from the models that ViperGPT used (see Section A.2).
@@
Strong baselines are essential for correctly measuring progress in machine learning. This is especially true in the emerging area of “tool use” (Cobbe et al., 2021; Komeili et al., 2021; Thoppilan et al., 2022; Parisi et al., 2022; Zeng et al., 2022; Gao et al., 2023; Qin et al., 2023; Zhuge et al., 2023). When using an LLM and other pre-trained models, we must be careful to report the exact LLM version and/or API when it was accessed, and ideally report results over several random seeds to measure the statistical significance. In Table 1, we provide an overview of the previous Zero-Shot (ZS), Few-Shot (FS), and Supervised (Sup) SotA methods, ViperGPT, end-to-end (E2E) baselines, and our results on all datasets we used for evaluation.
Early in the project, we found it difficult to reproduce the results reported in the ViperGPT paper. Our first hypothesis is that this is due to differences in the vision and language models we use compared to the ViperGPT paper. However, when running the original ViperGPT code from the official GitHub repository on RefCOCO, we were only able to achieve an IoU of 60.5 as opposed to 72.0 reported in the paper. Note, however, that ViperGPT uses Codex, which is discontinued, so we use GPT-3.5-turbo (OpenAI, 2023). Also note that this score was obtained using 16 dataset-specific ICEs (DS-ICEs). These examples contain large amounts of dataset-specific human-engineered information, such as “clothing requires returning the person”. In the case of truly Zero-Shot learning (without any human-engineered ICEs), the IoU score of ViperGPT’s official GitHub code drops by 14 points to 46.7. Moreover, in their GitHub code we found hand-engineered improvements: if returning a bounding box fails, then return an “average” bounding box, and if code execution fails on GQA, then query the image captioner (BLIP-2 (Li et al., 2023)). These code changes lead to improved results, but make it hard to quantify the true power of LLMs as controllers approach.
@@ 4 Related work
Recently, SotA on VQA has been largely obtained by scaling end-to-end vision and language models (VLMs) in terms of their size, training data, and compute (Alayrac et al., 2022; Chen et al., 2022b; Yu et al., 2022; Wang et al., 2022a; Gan et al., 2022; Lu et al., 2022; Li et al., 2023; Driess et al., 2023; Chen et al., 2023d; c). One of the earliest VLMs Flamingo (Alayrac et al., 2022) used a frozen pretrained language model of up to 70B parameters and a frozen pretrained image encoder with 435M parameters and trained only a “cross-attention” module that served as an interface between them. Since then, efforts have mainly gone into scaling both the image and the language models: GIT (Wang et al., 2022a) used a 300M language and scaled the image encoder to 4.8B parameters; PaLI (Chen et al., 2022b) scaled both components jointly, language model to 17B and image encoder to 4B parameters; PaLI-X (Chen et al., 2023d) continued this trend of scaling the total number of parameters to 55B by using an image encoder with 22B parameters; PaLM-E scaled the number of total parameters in the VLM to 562B by integrating the 540B PaLM (Chowdhery et al., 2022) and the 22B Vision Transformer (Dosovitskiy et al., 2020; Dehghani et al., 2023). On the other hand, BLIP-2 (Li et al., 2023) achieved SotA performance on various tasks with a 12B VLM and PaLI-3 Chen et al. (2023d) introduced a significantly smaller VLM with 5B total parameters that achieves competitive performance with SotA models on various VLM benchmarks. These models are typically pretrained on large amounts of data and then fine-tuned for the best performance on the downstream tasks. In contrast to this, visual reasoning with LLMs as programmers (Gupta & Kembhavi, 2023; Surís et al., 2023; Subramanian et al., 2023) does not require any fine-tuning or gradient updates on the downstream tasks. Moreover, even the largest VLMs struggle on the tasks that require compositional reasoning, the ability to generalize, fine-grained spatial capabilities, and counting (Bugliarello et al., 2023; Hsieh et al., 2023; Tschannen et al., 2023). Further scaling makes them even more data- and compute-hungry; therefore, it is unclear whether scaling alone can solve these tasks. Conversely, using LLMs as programmers enables task decomposition into subtasks and holds promise of strong generalization and compositional reasoning.
@@
Neural Modular Networks (NMNs) (Andreas et al., 2016; Johnson et al., 2017; Hu et al., 2017) are an alternative approach to monolithic end-to-end networks and offer a potential route to (compositional) generalization. They are also typically trained end-to-end, using supervised or reinforcement learning. NMNs are designed to have several modules, and the hope is that during training, each module will learn a different functionality, which will then be reusable across tasks. However, these models have a number of drawbacks: the program generation requires hand-tuned parsers, and they require optimization through reinforcement learning (e.g., REINFORCE Williams (1992)), which is often unstable. Learning all modules end-to-end hinders their ability to generalize (Bahdanau et al., 2018) and sometimes leads to a mode “collapse”, where some modules take over all the work and other modules are never activated, or modules that do not learn intended functionalities (Subramanian et al., 2020). Furthermore, they sometimes require supervision for program learning, which is difficult to obtain at scale. On the other hand, visual reasoning with LLMs as programmers mitigates many issues of NMNs: it does not require gradient-based training or finetuning, it is able to incorporate any modules or swap the existing ones, it leverages the strong generalization ability of LLMs, and it generates programs by utilizing the in-context learning ability of LLMs, thereby removing the need for training program generators. The programs generated by LLM do not have to be domain specific, but can use a common language such as Python that does not require custom interpreters.
@@
The field of LLMs as controllers for visual reasoning has received a great deal of interest recently. LLMs as controllers (also known as “tool use”) approach became prominent in the literature (Parisi et al., 2022; Schick et al., 2023), in particular for structured reasoning in the natural language domain (Madaan et al., 2022; Wang et al., 2023b; Gao et al., 2023; Chen et al., 2022a). In the domain of using LLMs as controllers for visual reasoning, PICa (Yang et al., 2022) solves a knowledge-based VQA task by first extracting an object and captions from the image and then querying GPT-3 with this information and in-context examples to answer a question. Socratic models (Zeng et al., 2022), HuggingGPT (Shen et al., 2023), and Societies of Mind (Zhuge et al., 2023) compose vision and language models to “communicate” in a fixed “protocol” and solve tasks such as image captioning, visual question answering, image generation, and robot planning. On the other hand, models such as VisProg (Gupta & Kembhavi, 2023), ViperGPT (Surís et al., 2023) and CodeVQA (Subramanian et al., 2023) go beyond a fixed communication “protocol” by having LLM write (Python) programs. During execution, the program calls individual vision modules (such as the object detector and depth estimator) through an API that is provided in the prompt. Additionally, VisProg is also capable of generating images as program output. These models showed great performance and achieved SotA on tasks such as compositional visual question answering, visual grounding, and video temporal reasoning. However, in their current form, these models rely on heavy human engineering of query-code examples in the prompt that are dataset- and task-specific and require significant labor by highly skilled workers. Our framework, on the other hand, is able to automatically generate in-context examples, removing the need for humans to write query-code examples, uses an Abstract API that puts less burden on the LLM to have strong spatial and temporal reasoning abilities, and proposes a way to “self-correct” the programs that failed to execute. Note that our framework is equally applicable to all of the above approaches.
@@
Vast literature shows that prompt format and contents are often important for achieving good performance with an LLM (Reynolds & McDonell, 2021; Zhao et al., 2021; Lu et al., 2021; Moradi & Samwald, 2021; Madaan & Yazdanbakhsh, 2022; Wei et al., 2023). A prompt typically consists of a task description (in natural language), in-context examples (e.g. query-code in ViperGPT) and an API (in the case where LLMs write a program). Various prompting techniques have been engineered, such as Chain-of-Thought prompting (Wei et al., 2022), Self-Consistency (Wang et al., 2022c), Tree of Thoughts (Yao et al., 2023), Graph of Thoughts (Besta et al., 2023), Plan-and-Solve Prompting (Wang et al., 2023a), Least-to-Most Prompting (Zhou et al., 2022a), etc. All these techniques rely on human prompt engineering, in particular, on in-context examples. On the other hand, some methods try to automate prompt engineering. They sometimes use gradient-based optimization (Shin et al., 2020; Gao et al., 2020; Wen et al., 2023) and some approaches require only API access to the model (Xu et al., 2022; Prasad et al., 2022). Other works use LLMs for prompt optimization. APE (Zhou et al., 2022b) first generates instructions with an LLM, then selects instructions with the highest accuracy, and uses them for future LLM prompts. APO (Pryzant et al., 2023) generates feedback with an LLM that informs how to update the previous instruction. OPRO (Yang et al., 2023a) uses an LLM to generate new instructions at each optimization step, asking the LLM to improve task accuracy by changing task instructions, which requires determining the score on a small set of labeled examples and providing it in the meta-prompt. Promptbreeder (Fernando et al., 2023) goes a step further and proposes a self-referential self-improvement LLM using a meta-prompt that controls the generation of the main (task) prompt and evolves both via mutation. More importantly, Promptbreeder shows some surprising results such that a simple prompt “SOLUTION” outperforms all previous approaches. This further demonstrates the sensitivity of LLMs and the importance of automatizing the prompt engineering process. Common to all above frameworks for automatizing prompting is that they automatize the “task description” part of the prompt. On the other hand, in our framework, we automatize the generation of in-context examples, which might have an even greater influence on the performance of the LLM.
@@
In the LLM literature, there have been mixed findings on the ability of LLMs to critique and self-correct their own reasoning and outputs. Self-Refine (Madaan et al., 2023) provides feedback to the LLM of the previously generated output, which is then refined. Several other approaches show benefits of providing feedback to LLM in improving reasoning (Shinn et al., 2023; Madaan et al., 2023), code generation (Chen et al., 2023e; Olausson et al., 2023; Chen et al., 2023a), improving LLM alignment (Bai et al., 2022; Ganguli et al., 2023), etc. On the other hand, there has been increasing evidence that LLMs cannot self-correct reasoning yet (Huang et al., 2023; Stechly et al., 2023; Valmeekam et al., 2023), unless they receive external feedback, which usually requires access to ground truth labels. In our work, we also found that providing the previous question and code as feedback to the model did not improve the results. However, we show that it is possible to improve performance by tuning hyperparameters on the fly, a direction that, to the best of our knowledge, has not been explored previously.

### v2

@@ 1 Introduction
Compositional visual question answering requires a model to answer questions about visual content in a compositional manner, involving multiple steps of reasoning or considering relationships between different objects or entities within an image. It is a complex task as it requires understanding both the visual information in an image and the structure of the question, and reasoning about how different visual elements relate to one another to generate the correct answer. Recently, large progress has been made on many such vision and language tasks by scaling end-to-end neural networks models in terms of size, training data, and compute (Alayrac et al., 2022; Chen et al., 2023c; Yu et al., 2023; Wang et al., 2022a; Gan et al., 2022; Lu et al., 2023; Li et al., 2023b; Driess et al., 2023; Chen et al., 2023b; 2024a). However, even the largest state-of-the-art (SotA) models struggle in tasks that require compositional reasoning, ability to generalize, fine-grained spatial reasoning capabilities, and counting (Bugliarello et al., 2023; Paiss et al., 2023; Hsieh et al., 2023; Yuksekgonul et al., 2022; Zhao et al., 2022; Hendricks & Nematzadeh, 2021). An example of such task is the following query: “Could the cookies on the table be equally distributed among children?” (Surís et al., 2023). To solve this, the model needs to detect the cookies in the image, filter out the ones that are not on the table, detect children, count cookies and children, and check if the cookies count is divisible by the children count. Questions like these are difficult for current end-to-end vision and language models (VLMs). Scaling VLMs further makes them even more data- and compute-hungry(Villalobos et al., 2022), so the scale alone seems unlikely to solve these tasks, especially due to the exponentially-large long tail of compositional tasks.
@@
Recently, an alternative approach based on “tool use” gained popularity (Cobbe et al., 2021; Komeili et al., 2022; Thoppilan et al., 2022; Parisi et al., 2022; Zeng et al., 2022a; Gao et al., 2023a; Qin et al., 2024; Zhuge et al., 2023). In “tool use”, an LLM solves a task by controlling (akin to System 2) a set of tools (such as an object detector, akin to System 1) (Zeng et al., 2022a; Shen et al., 2023; Zhuge et al., 2023). In particular, VisProg (Gupta & Kembhavi, 2023), ViperGPT (Surís et al., 2023), and CodeVQA (Subramanian et al., 2023) show great promise in solving visual question answering by using an LLM to generate a program (in Python or a custom scripting language). During execution, the program calls individual vision models (such as object detector, depth estimator) through an API that is provided in the prompt. For example, to answer “Which color is the jacket of the second person from the left?” (Figure 1(a)), the program needs to detect people, sort them from left to right, select the second, detect their jacket, and query its color. These models achieved SotA on compositional visual question answering, visual grounding, and video temporal reasoning tasks. By their construction, they are interpretable, compositional, adaptable (tools can be upgraded on the fly), offer strong generalization, mathematical, and reasoning skills, and do not require gradient-based training. However, in their current form, they heavily rely on human engineering of in-context (program) examples (ICEs) in the prompt. Moreover, ICEs are dataset- and task-specific. To generate them, significant labor by highly skilled programmers is required. For this reason, we argue that these methods should not be called zero-shot in their current form.
@@ 2.2 Abstract API through visual routines
When programming, we continuously build new layers of abstraction. grouping them together into new functions. By abstracting away the implementation details, we reduce the cognitive load on the programmer which is then able to build systems of increased complexity. Motivated by this, we introduce a set of spatially and temporally abstract functions (routines  [^1] Note that our “routines” do not correspond to the visual routines of Ullman (1987) such as tracing or scanning.) that abstract away a number of lines of code for the same functionality and together make the Abstract API. From a practical perspective, we are motivated by a number of failure cases observed in the experiments (see Section 3). As presented in Section 2.1, ViperGPT’s API is fairly simple (contains almost exclusively functions to call pretrained models). Although simplicity is good and often desirable, in the case of visual reasoning with LLMs as programmers, the lack of expressive visual routines requires the code-generating LLM to have strong spatial and temporal reasoning capabilities. Qualitative analysis showed that this is often not the case and that the current LLMs are not yet perfect in these terms (e.g. they confuse left and right, top and bottom). For example, for the query “return the second person from the right”, the program generated by the LLM correctly sorts the persons along the horizontal axis but then wrongly takes the second index in the array (instead of the second last). Similarly, they sometimes “confuse” temporal order, e.g., if a “before” event means a smaller or a larger time index.
@@ 2.3 Automatic generation of in-context examples
In-context examples (ICEs) greatly influence the performance of LLMs (Brown et al., 2020; Chen et al., 2023). For example, ViperGPT (Surís et al., 2023) uses between 10 and 16 hand-engineered dataset-specific query-code ICEs per dataset. Similarly, VisProg (Gupta & Kembhavi, 2023) uses between 16 and 31 ICEs and CodeVQA (Subramanian et al., 2023) about 50 ICEs. However, constructing these ICEs requires heavy human engineering, as they might need to be rewritten in a way that the LLM can use them to “generalize” to other examples across the dataset. Furthermore, the constructed examples are specific not only to the dataset but also to the LLM and the API. If any of those changes, they need to be written from scratch. Finally, to write good query-code ICEs, highly skilled labor is required, ideally someone familiar with the workings of LLMs and a good grasp of Python programming.
@@ 2.4 Self-correction
In some cases, we know that code execution failed due to some components of a particular module. For example, the open vocabulary detector fails due to a too high threshold hyperparameter. When the threshold is high, we have a higher number of false negatives. For such cases, we propose to automatically change the hyperparameter of the module (e.g. reduce the threshold) and execute code again. Although here we experiment with only a single option of tuning such hyperparameters, the idea is applicable to any model that involves such “sensitivity” hyperparameters. For example, another such candidate could be the threshold that determines the “similarity” of the CLIP-style image-text embedding model. The task of this module is to filter out a particular instance (e.g. “blue ball”) from a set of instances (“balls”), e.g., based on an attribute (“blue”). It does so by comparing the text embedding of “blue” with the image embeddings of all “ball” image patches in the image. Then it returns all image patches whose embeddings have similarity with the text embedding higher than the threshold value. Here we would similarly tune the threshold in case this module fails.
@@ 3 Experiments
For code generation, we use a code instruction-tuned version of PaLM 2 (Anil et al., 2023) code-bison accessible via the Google Cloud API (Google, 2023). We use the same model to select an answer for the multichoice questions in the NExT-QA dataset. Vision models we use are OWLv2 (Minderer et al., 2023) for object detection, SigLiT (Zhai et al., 2023) for text-image comparison, MiDaS (Ranftl et al., 2020) for depth estimation, and PaLI-3 (Chen et al., 2023b) for image captioning and answering visual queries. Note that all models are different from the models that ViperGPT used (see Section A.2).
@@
Strong baselines are essential for correctly measuring progress in machine learning. This is especially true in the emerging area of “tool use” (Cobbe et al., 2021; Komeili et al., 2022; Thoppilan et al., 2022; Parisi et al., 2022; Zeng et al., 2022a; Gao et al., 2023a; Qin et al., 2024; Zhuge et al., 2023). When using an LLM and other pre-trained models, we must be careful to report the exact LLM version and/or API when it was accessed, and ideally report results over several random seeds to measure the statistical significance. In Table 1, we provide an overview of the previous Zero-Shot (ZS), Few-Shot (FS), and Supervised (Sup) SotA methods, ViperGPT, end-to-end (E2E) baselines, and our results on all datasets we used for evaluation.
Early in the project, we found it difficult to reproduce the results reported in the ViperGPT paper. Our first hypothesis is that this is due to differences in the vision and language models we use compared to the ViperGPT paper. However, when running the original ViperGPT code from the official GitHub repository on RefCOCO, we were only able to achieve an IoU of 60.5 as opposed to 72.0 reported in the paper. Note, however, that ViperGPT uses Codex, which is discontinued, so we use GPT-3.5-turbo (OpenAI, 2023). Also note that this score was obtained using 16 dataset-specific ICEs (DS-ICEs). These examples contain large amounts of dataset-specific human-engineered information, such as “clothing requires returning the person”. In the case of truly Zero-Shot learning (without any human-engineered ICEs), the IoU score of ViperGPT’s official GitHub code drops by 14 points to 46.7. Moreover, in their GitHub code we found hand-engineered improvements: if returning a bounding box fails, then return an “average” bounding box, and if code execution fails on GQA, then query the image captioner (BLIP-2 (Li et al., 2023b)). These code changes lead to improved results, but make it hard to quantify the true power of LLMs as controllers approach.
@@ 4 Related work
Recently, SotA on VQA has been largely obtained by scaling end-to-end vision and language models (VLMs) in terms of their size, training data, and compute (Alayrac et al., 2022; Chen et al., 2023c; Yu et al., 2023; Wang et al., 2022a; Gan et al., 2022; Lu et al., 2023; Li et al., 2023b; Driess et al., 2023; Chen et al., 2023b; 2024a). One of the earliest VLMs Flamingo (Alayrac et al., 2022) used a frozen pretrained language model of up to 70B parameters and a frozen pretrained image encoder with 435M parameters and trained only a “cross-attention” module that served as an interface between them. Since then, efforts have mainly gone into scaling both the image and the language models: GIT (Wang et al., 2022a) used a 300M language and scaled the image encoder to 4.8B parameters; PaLI (Chen et al., 2023c) scaled both components jointly, language model to 17B and image encoder to 4B parameters; PaLI-X (Chen et al., 2023b) continued this trend of scaling the total number of parameters to 55B by using an image encoder with 22B parameters; PaLM-E scaled the number of total parameters in the VLM to 562B by integrating the 540B PaLM (Chowdhery et al., 2023) and the 22B Vision Transformer (Dosovitskiy et al., 2021; Dehghani et al., 2023). On the other hand, BLIP-2 (Li et al., 2023b) achieved SotA performance on various tasks with a 12B VLM and PaLI-3 Chen et al. (2023b) introduced a significantly smaller VLM with 5B total parameters that achieves competitive performance with SotA models on various VLM benchmarks. These models are typically pretrained on large amounts of data and then fine-tuned for the best performance on the downstream tasks. In contrast to this, visual reasoning with LLMs as programmers (Gupta & Kembhavi, 2023; Surís et al., 2023; Subramanian et al., 2023) does not require any fine-tuning or gradient updates on the downstream tasks. Moreover, even the largest VLMs struggle on the tasks that require compositional reasoning, the ability to generalize, fine-grained spatial capabilities, and counting (Bugliarello et al., 2023; Hsieh et al., 2023; Tschannen et al., 2023). Further scaling makes them even more data- and compute-hungry; therefore, it is unclear whether scaling alone can solve these tasks. Conversely, using LLMs as programmers enables task decomposition into subtasks and holds promise of strong generalization and compositional reasoning.
@@
Neural Modular Networks (NMNs) (Andreas et al., 2016; Johnson et al., 2017; Hu et al., 2017) are an alternative approach to monolithic end-to-end networks and offer a potential route to (compositional) generalization. They are also typically trained end-to-end, using supervised or reinforcement learning. NMNs are designed to have several modules, and the hope is that during training, each module will learn a different functionality, which will then be reusable across tasks. However, these models have a number of drawbacks: the program generation requires hand-tuned parsers, and they require optimization through reinforcement learning (e.g., REINFORCE Williams (1992)), which is often unstable. Learning all modules end-to-end hinders their ability to generalize (Bahdanau et al., 2019) and sometimes leads to a mode “collapse”, where some modules take over all the work and other modules are never activated, or modules that do not learn intended functionalities (Subramanian et al., 2020). Furthermore, they sometimes require supervision for program learning, which is difficult to obtain at scale. On the other hand, visual reasoning with LLMs as programmers mitigates many issues of NMNs: it does not require gradient-based training or finetuning, it is able to incorporate any modules or swap the existing ones, it leverages the strong generalization ability of LLMs, and it generates programs by utilizing the in-context learning ability of LLMs, thereby removing the need for training program generators. The programs generated by LLM do not have to be domain specific, but can use a common language such as Python that does not require custom interpreters.
@@
The field of LLMs as controllers for visual reasoning has received a great deal of interest recently. LLMs as controllers (also known as “tool use”) approach became prominent in the literature (Parisi et al., 2022; Schick et al., 2023), in particular for structured reasoning in the natural language domain (Madaan et al., 2022; Wang et al., 2023b; Gao et al., 2023a; Chen et al., 2023a). In the domain of using LLMs as controllers for visual reasoning, PICa (Yang et al., 2022) solves a knowledge-based VQA task by first extracting an object and captions from the image and then querying GPT-3 with this information and in-context examples to answer a question. Socratic models (Zeng et al., 2022a), HuggingGPT (Shen et al., 2023), Societies of Mind (Zhuge et al., 2023), and many other papers (Liu et al., 2023; Zhu et al., 2024b; Zhang et al., 2024; Li et al., 2023c; Yang et al., 2024b; Xu et al., 2023; Zhang et al., 2023; Gao et al., 2023b; Ye et al., 2023b; Yang et al., 2023b; Lu et al., 2024; Wu et al., 2023; Zhu et al., 2024a) (for a comprehensive survey, see Yin et al. (2023) or Li et al. (2023a)) compose vision and language models to “communicate” in a fixed “protocol” and solve tasks such as image captioning, visual question answering, image generation, and robot planning. On the other hand, models such as VisProg (Gupta & Kembhavi, 2023), ViperGPT (Surís et al., 2023) and CodeVQA (Subramanian et al., 2023) go beyond a fixed communication “protocol” by having LLM write (Python) programs. During execution, the program calls individual vision modules (such as the object detector and depth estimator) through an API that is provided in the prompt. Additionally, VisProg is also capable of generating images as program output. These models showed great performance and achieved SotA on tasks such as compositional visual question answering, visual grounding, and video temporal reasoning. However, in their current form, these models rely on heavy human engineering of query-code examples in the prompt that are dataset- and task-specific and require significant labor by highly skilled workers. Our framework, on the other hand, is able to automatically generate in-context examples, removing the need for humans to write query-code examples, uses an Abstract API that puts less burden on the LLM to have strong spatial and temporal reasoning abilities, and proposes a way to “self-correct” the programs that failed to execute. Note that our framework is equally applicable to all of the above approaches.
@@
Vast literature shows that prompt format and contents are often important for achieving good performance with an LLM (Reynolds & McDonell, 2021; Zhao et al., 2021; Lu et al., 2022; Moradi & Samwald, 2021; Madaan & Yazdanbakhsh, 2022; Wei et al., 2023). A prompt typically consists of a task description (in natural language), in-context examples (e.g. query-code in ViperGPT) and an API (in the case where LLMs write a program). Various prompting techniques have been engineered, such as Chain-of-Thought prompting (Wei et al., 2022), Self-Consistency (Wang et al., 2023c), Tree of Thoughts (Yao et al., 2023), Graph of Thoughts (Besta et al., 2024), Plan-and-Solve Prompting (Wang et al., 2023a), Least-to-Most Prompting (Zhou et al., 2023a), etc. All these techniques rely on human prompt engineering, in particular, on in-context examples. On the other hand, some methods try to automate prompt engineering. They sometimes use gradient-based optimization (Shin et al., 2020; Gao et al., 2021; Wen et al., 2023) and some approaches require only API access to the model (Xu et al., 2022; Prasad et al., 2023). Other works use LLMs for prompt optimization. APE (Zhou et al., 2023b) first generates instructions with an LLM, then selects instructions with the highest accuracy, and uses them for future LLM prompts. APO (Pryzant et al., 2023) generates feedback with an LLM that informs how to update the previous instruction. OPRO (Yang et al., 2024a) uses an LLM to generate new instructions at each optimization step, asking the LLM to improve task accuracy by changing task instructions, which requires determining the score on a small set of labeled examples and providing it in the meta-prompt. Promptbreeder (Fernando et al., 2023) goes a step further and proposes a self-referential self-improvement LLM using a meta-prompt that controls the generation of the main (task) prompt and evolves both via mutation. More importantly, Promptbreeder shows some surprising results such that a simple prompt “SOLUTION” outperforms all previous approaches. This further demonstrates the sensitivity of LLMs and the importance of automatizing the prompt engineering process. Common to all above frameworks for automatizing prompting is that they automatize the “task description” part of the prompt. On the other hand, in our framework, we automatize the generation of in-context examples, which might have an even greater influence on the performance of the LLM.
@@
In the LLM literature, there have been mixed findings on the ability of LLMs to critique and self-correct their own reasoning and outputs. Self-Refine (Madaan et al., 2023) provides feedback to the LLM of the previously generated output, which is then refined. Several other approaches show benefits of providing feedback to LLM in improving reasoning (Shinn et al., 2024; Madaan et al., 2023), code generation (Chen et al., 2024b; Olausson et al., 2023; Chen et al., 2023), improving LLM alignment (Bai et al., 2022; Ganguli et al., 2023), etc. On the other hand, there has been increasing evidence that LLMs cannot self-correct reasoning yet (Huang et al., 2024; Stechly et al., 2023; Valmeekam et al., 2023), unless they receive external feedback, which usually requires access to ground truth labels. In our work, we also found that providing the previous question and code as feedback to the model did not improve the results. However, we show that it is possible to improve performance by tuning hyperparameters on the fly, a direction that, to the best of our knowledge, has not been explored previously.